OpenStudy (anonymous):

Solve the given equation. Check your solution. -3|2n + 5| = -9

OpenStudy (anonymous):

-9/-3 = |2n + 5| 3 = |2n + 5| 2n + 5 = 3 2n + 5 = -3 2n = 3 - 5 2n = -3 -5 2n = -2 2n = -8 n = -1 n = -4

OpenStudy (anonymous):

okay i see how to do it, thank you

OpenStudy (anonymous):

yw